OpenStudy (anonymous):

The ratio of the angle measures of the acute angles in a right triangle is 1:2

OpenStudy (anonymous):



OpenStudy (anonymous):

if a is a right angle, then b is bigger than c in a ratio of 2:1 180-90=90 there are 3 parts to the ratio (2+1) so we divide 90 by 3 90/3=30 b=2*30=60 c=1*30=30

OpenStudy (anonymous):

A right triangle is a triangle with a 90 degree angle, correct? Stemming from this, the other two have to add up to 90 degrees since a triangle, by definition, always has 180 degrees total in its interior angles. Therefore, to find the other two angles you may use the algebraic equation 3x=90 (since there are three measurements of x - angle 1 is one and angle two is twice that) So x= 30 - the triangle is a 30, 60, 90 with the side one being 30 and side two 60.

OpenStudy (anonymous):

The measure of one angle of a triangle is 40. The measures of the 2 angles are in a ratio of 3:4